OpenStudy (erniellyd28):

sarah took the advertising department from her company on a round trip to meet with a client. including sarah a total of 6 people took the trip.she was able to purchase coach tickets for $250 and first class tickets for $940.she used her total budget for airfare for the trip,which was $4260.how many first class tickets did she buy? how many coach tickets did she buy?

OpenStudy (plainntall):

It helps to write things out. First let's make up two variables Coach Tickets = C First Class Tickets = F Next write down the given relationships 6 people mean 6 tickets giving us C + F = 6 Cost is the number of First Class Tickets times $940 plus number of Coach tickets times $250 with the total equal to $4260 $940F+$250C= $4260 Now you have two equations solving C + F = 6 for either C or F will allow you to substitute it into the second equation giving you only one variable to solve for. Once you have that variable subtract it from 6 to get the second.

OpenStudy (erniellyd28):

so the answer is 6 for both?

OpenStudy (plainntall):

That would give 12 people on the trip. Solving C+F=6 for C gives C=6-F now substitute it into $940F + $250C = $940F +$250(6-F) = $4260 What does F equal?

OpenStudy (erniellyd28):

250(6-f)

OpenStudy (plainntall):

940F + 250(6-F)=4260 940F + 1500 - 250F = 4260 (multiplying 250x6 and 250xF) (940-250)F + 1500 = 4260 690F + 1500 = 4260 690F + 1500 -1500 = 4260 -1500 (subtract 1500 from both sides) 690F=2760 F = 2760/690 = 4 Now that we know F we can find C C=6-F = 6-4 = 2 4 first class tickets and 2 Coach

OpenStudy (plainntall):

1. read the problem 2. make variables 3. right down relationship in equation form 4. Solve for one variable 5. substitute for that solved variable 6. Solve for the second variable 5 Solve for the first variable using the second variable

OpenStudy (erniellyd28):

thank you very very much for your time and help :-)

OpenStudy (plainntall):

welcome

OpenStudy (erniellyd28):

sorry if i was difficult to work with lol

OpenStudy (plainntall):

No reason to apologize. I am helping because I want to. Math is challenging, but it can be fun once you figure it out and that can be difficult.